BOND VALUATION An investor has two bonds in her portfolio, Bond C and Bond Z. Each
bond matures in 4 years, has a face value of $1,000, and has a yield to maturity of 9 6%.
Bond C pays a 10% annual coupon, while Bond Z is a zero coupon bond.
a. Assuming that the yield to maturity of each bond remains at 9 6% over the next 4
years, calculate the price of the bonds at each of the following years to maturity:

          

Years to Maturity                  Bond C Price                                     Bond Z Price

                        4 _____________    _____________  

                        3    _____________     _____________  

                        2    _____________     _____________

                        1     _____________    _____________  

0     _____________     _____________  

b. Plot the time path of prices for each bond.- 7-6 BOND VALUATION An investor has two bonds in her portfolio, Bond C and Bond Z. Each bond matures in 4 years, has a face
Answer #1

ANS -

The bond C & Z value has been calculated by discounting each cash flow till year 0

Bond Z is zero coupon bond , so the price is is also low because there are no cash flow except last year cash flow

Bond C face volue = $100 Coupon - 104. on face rowe - 10. X loom = $ 100 YTM = 9.6% т Чувал ,CE = 100 CF₂ = 100 CF4 = 100+ 10Bond 2 (zew Coupon Bond) YTM = 9.6% pa T 4 F: 1000 Year = 1000 (110964 L = $693.03 OR Year 3 (ena) - 1000 (11096) - 912.40 Ye